# Consensus Planning with Primal, Dual, and Proximal Agents

### v2

@@ 1 Introduction
Consensus planning refers to a coordination mechanism to align different distributed agents (or systems) who share common resources and who must act in order to achieve a joint goal. For example, in the context of a large retailer, such a distribution of tasks is necessary, given the scale of the supply chain problem, with separate systems focusing on (say) buying, removals, placement, capacity control, fulfillment, or transfers, to name just a few. These individual components act semi-independently, but some form of communication is required, given their interdependence, to ensure that they work in unison. This alignment is often times built through ad-hoc solutions, or proxies of connected systems. To overcome these limitations, distributed, market-based approaches, such as a Consensus Planning Protocol (CPP) [DPJvR22, vR24], can be developed within organizations. A CPP starts with the agents, and it introduces a coordinator. Information is exchanged between the agents through the coordinator in a structured manner; and the coordinator reconciles the various favored plans of the individual agents in a principled way to eventually reach a consensus. Associated with such mechanisms are agent-level plans, costs that characterize their favored plans, and costs of deviating from them.
@@
Consensus planning is often viewed as a variant of distributed optimization, where there are relatively few optimization agents and querying information from each agent is comparatively costly. A CPP relies on the iterative exchange of information between agents and a coordinator. The mechanism through which the coordinator updates and aligns the plans and prices hinges upon the nature of the agents. A theoretically-popular approach centers around the use of proximal interfaces, which are motivated by the theoretically-attractive properties of the Alternative Direction Method of Multipliers (ADMM) [BPC+11]. In this proximal framework, the agents and the coordinator exchange two types of information: 1) a favored plan (primal variable); and 2) a price (or cost of deviation) (dual variable). These two pieces of information are iteratively exchanged and updated by the agents and coordinator until convergence. Implicit in this approach of using ADMM as a coordination mechanism is the assumption that all agents are amenable to interacting with the coordinator via a proximal interface.
@@
More generally, we can categorize agents into three broad classes, depending on the interface that they support: primal, dual, and proximal. Primal agents can be probed given a tentative plan, and they return a corresponding cost of deviation; dual agents can be called with a price, and they return a corresponding favored plan; and proximal agent can consume both a tentative plan and tentative cost, and they return updated values thereof. Most CPP algorithms are built with one type of agents in mind, requiring that all agents belong to the same class. The practical difficulty that we address theoretically in this paper is to devise methodologies that allow for a “mix-and-match” of agents, that are compatible with all types of agent interfaces, so as not to require the preliminary burden of modifying them to fit the same type.
@@ 2.1 Consensus Problem
The consensus problem is one that involves a set of agents $\mathcal{M}$ that each have their own cost function $g_{i}(x_{i}),i\in\mathcal{M}$. The goal is to optimize the sum of these functions for a common plan $z$: $\min_{z}\sum_{i\in\mathcal{M}}g_{i}(z)$. In order to leverage the separability of the agents, the problem is re-written to endow each agent with their own plan variable, subject to the constraint that they must all be equal:
$\min_{x_{i},z}$ $\leavevmode\nobreak\ \sum_{i\in\mathcal{M}}g_{i}(x_{i})$
$s.t.$ $\leavevmode\nobreak\ x_{i}=z,\leavevmode\nobreak\ \forall i\in\mathcal{M}.$
@@ 2.2 Agents
In our consensus planning problem, there exist three types of agent interfaces, primal, dual, and proximal, which are the natural interfaces for the three main approaches to the problem, gradient descent, dual ascent, and the augmented Lagrangian method of multipliers, respectively.
@@
$\max_{\lambda_{i}}\left\{\min_{x_{i}}\sum_{i\in\mathcal{M}}g_{i}(x_{i})-\lambda_{i}^{T}(z-x_{i})\right\}.$
We then require from the agents that they be able to consume the dual variables $\lambda_{i}$ and return the solutions to the subproblems in bracket as $x_{i}=\arg\min_{x}g_{i}(x)-\lambda_{i}^{T}(z-x)$.
@@
$\max_{\lambda_{i}}\left\{\min_{x_{i}}\sum_{i\in\mathcal{M}}g_{i}(x_{i})-\lambda_{i}^{T}(z-x_{i})+\frac{\rho}{2}\lVert z-x_{i}\rVert^{2}\right\}.$
@@ 2.3 Previous Work
Consensus planning is often viewed as a variant of distributed optimization, where there are relatively few optimization agents and querying information from each agent is comparatively costly. This variant is in contrast with much of the distributed optimization work where the central concern is that of scale across many agents [RY22, ch.11]. Most of the distributed optimization algorithms assume the same interface for all agents: for example, a primal interface in the case of distributed (proximal) gradient descent; a dual interface in the case of dual decomposition; and a proximal interface in the case of distributed ADMM. By contrast, this paper allows for any mix of primal, dual, and proximal interfaces.
@@ 2.4 Motivating Examples
Table 1: Examples of consensus problems at a large online retailer.
@@ 3.1 Derivation
We consider a consensus optimization problem involving a mixed set of agents $i\in\mathcal{M}$ that can be either primal, dual, or proximal. We let $\mathcal{P}$ be the set of primal agents, $\mathcal{D}$ be the set of dual agents, and $\mathcal{X}$ be the set of proximal agents. Each agent’s objective function is given by $g_{i}$, where $i$ denotes their index, and we assume that the functions $g_{i}$ are convex, and in particular $\mu_{i}$-strictly convex for $i\in\mathcal{D}$, and have $\beta_{i}$-Lipschitz continuous gradients for $i\in\mathcal{X}$. The consensus problem to be solved is:
$\min_{\mathbf{x},z}$ $\leavevmode\nobreak\ \sum_{i\in\mathcal{M}}g_{i}(x_{i})$ (1)
@@
We first make the different treatments of the agents explicit by separating agents based on their type. Applying partial duality (e.g., [LY21, ch. 14]) to the dual and proximal agents, and augmenting the latter with the augmented Lagrangian proximal term associated with their corresponding constraint, yields the following equivalent formulation:
$\max_{\lambda_{i},i\in\mathcal{D}\cup\mathcal{X}}\min_{x_{i},z}$ $\leavevmode\nobreak\ \sum_{i\in\mathcal{P}}g_{i}(x_{i})+\sum_{i\in\mathcal{D}}g_{i}(x_{i})+\lambda_{i}^{T}(z-x_{i})+\sum_{i\in\mathcal{X}}g_{i}(x_{i})+\lambda_{i}^{T}(z-x_{i})+\frac{\rho_{i}}{2}\lVert z-x_{i}\rVert^{2},$ (2)
@@
Formulation (2) already suggests a direction for the solution of the problem, owing in particular to the similarity in structure between ADMM [BPC+11] and traditional dual ascent methods [LY21, ch. 14] used to solve dualized consensus problems. Both of these approaches involve iterations comprised of the same 3 steps:
@@
A natural idea is to recast the primal agent as a proximal agent, where its function $g_{i}$ is replaced by some linearized approximation thereof using the gradient information returned by the primal agent. Fortunately, a substantive body of work has demonstrated that the ADMM updates do not need to be exact for the method to converge [HLHY02, EB92, YZ11]. In particular, the terms of the ADMM updates can be linearized, be it the augmented quadratic term [LLS11, YY13], or the agent function [OHTG13, Suz13], or both through a more general use of a Bregman divergence term to replace the quadratic penalty, and an additional Bregman divergence term [WB14]. Leveraging these latter results, a simple Bregman-ADMM formulation in the context of the consensus problem has updates of the form:
@@
We will see in Section 3.3 that when all agents are primal, the application of this linearized ADMM yields updates that bear a strong similarity to the ones performed by distributed gradient descent [NO09].
@@ 3.2 Formulation
We detail in this section the steps of the algorithm following the high level derivation presented in Section 3.1. The agents can be either primal ($i\in\mathcal{P}$), dual ($i\in\mathcal{D}$), or proximal ($i\in\mathcal{X}$). We recall that we consider $\mu_{i}$-strongly convex functions $g_{i},i\in\mathcal{D}$ for dual agents, and $\beta_{i}$-Lipschitz continuous gradients for $g_{i},i\in\mathcal{P}$. We allow for $\mu_{i}=0$ and $L_{i}=+\infty$ for those other functions that are not strongly convex or do not have Lipschitz continuous gradients, for notational simplicity. Each agent is allowed to have their own regularizing parameter/learning rate $\rho_{i}$, although in practice we often use the same one for agents of the same type. We will return to a discussion on the choice of these hyperparameters in Section 5, but since the dual agents essentially perform dual ascent, and as we will confirm in the proof of the convergence results, we impose that $\rho_{i}<\mu_{i},\leavevmode\nobreak\ \in\mathcal{D}$. On the other hand, we recall that we require $L_{i}>\beta_{i}$ for $i\in\mathcal{P}$, and place no restriction on $\rho_{i}>0$ for $i\in\mathcal{X}$. We further assume throughout the paper that the initial prices $\lambda_{i}^{0}$ are such that $\sum_{i\in\mathcal{M}}\lambda_{i}=0$.
@@
$\phi_{i}(x)$ =$=\begin{cases}0,&i\in\mathcal{X}\cup\mathcal{D}\\ \frac{L_{i}}{2}\lVert x\rVert^{2}-g_{i}(x),&i\in\mathcal{\mathcal{P}}\end{cases},$ $\tilde{\rho}_{i}$ =$=\begin{cases}\rho_{i},&i\in\mathcal{\mathcal{P}\cup\mathcal{X}}\\ 0,&i\in\mathcal{D}.\end{cases}$
Note that, technically, the update of the primal agents is performed by the coordinator. The primal agents are queried with the tentative plan $x_{i}^{k}$ and return $\nabla g_{i}(x_{i}^{k})$, which is then used by the coordinator to yield the updated primal agent plan through (4). Additionally, the consensus update should have a term comprised of a multiple of the sum of the prices, but summing up the price updates, (8) we obtain the optimality condition of the consensus update problem, meaning that the sum of the prices is null after the first iteration (see e.g. [BPC+11]). Making the additional assumption that the sum of the initial prices is also null further simplifies the notation. The algorithm is summarized in Algorithm 1, which we name 3-Agent Consensus Planning (3ACP).
@@
Let $\rho_{i}>0,\leavevmode\nobreak\ \forall i\in\mathcal{M}$, and $\lambda_{i}^{0}$ such that $\sum_{i\in\mathcal{M}}\lambda_{i}^{0}=0$.
@@ 3.3 Comments
Consider the case when all the agents are primal agents, and assume for simplicity that they use common values $L=L_{i},\forall i$ and $\rho=\rho_{i},\forall i$. The updates take the form of Bregman ADMM [WB14]:
@@
Given that the matrix $W$ is symmetric and doubly stochastic, these updates are almost those of the decentralized gradient descent [NO09], except for the presence of dual variables $\lambda^{k}$ in the update. Decentralized gradient descent requires that the step sizes be decreasing in order to converge to the optimal, otherwise it only converges to within a neighborhood of the optimal [NO09, YLY16]. When the functions are strongly convex, the consensus plan converges linearly to the optimal until it reaches said neighborhood [YLY16]. By working on both primal and dual variables through the coordinator, the linearized (Bregman) ADMM overcomes the need for decreasing step sizes and converges linearly to the optimal solution.
@@
When all the agents are proximal, the algorithm reduces to the traditional consensus ADMM [BPC+11].

### v3

@@ 1 Introduction
Consensus planning refers to a coordination mechanism to align different distributed agents (or systems) who share common resources and who must act in order to achieve a joint goal. For example, in the context of a large retailer, such a distribution of tasks is necessary, given the scale of the supply chain problem, with separate systems focusing on (say) buying, removals, placement, capacity control, fulfillment, or transfers, to name just a few. These individual components act semi-independently, but some form of communication is required, given their interdependence, to ensure that they work in unison. This alignment is often times built through ad-hoc solutions, or proxies of connected systems. To overcome these limitations, distributed, market-based approaches, such as a Consensus Planning Protocol (CPP) (cpp; cpptalk), can be developed within organizations. A CPP starts with the agents, and it introduces a coordinator. Information is exchanged between the agents through the coordinator in a structured manner; and the coordinator reconciles the various favored plans of the individual agents in a principled way to eventually reach a consensus. Associated with such mechanisms are agent-level plans, costs that characterize their favored plans, and costs of deviating from them.
@@
Consensus planning is often viewed as a variant of distributed optimization, where there are relatively few optimization agents and querying information from each agent is comparatively costly. A CPP relies on the iterative exchange of information between agents and a coordinator. The mechanism through which the coordinator updates and aligns the plans and prices hinges upon the nature of the agents. A theoretically popular approach centers around the use of proximal interfaces, which are motivated by the theoretically attractive properties of the Alternative Direction Method of Multipliers (ADMM) (boyd2011distributed). In this proximal framework, the agents and the coordinator exchange two types of information: 1) a favored plan (primal variable); and 2) a price (or cost of deviation) (dual variable). These two pieces of information are iteratively exchanged and updated by the agents and coordinator until convergence. Implicit in this approach of using ADMM as a coordination mechanism is the assumption that all agents are amenable to interacting with the coordinator via a proximal interface.
@@
More generally, we can categorize agents into three broad classes, depending on the interface that they support: primal, dual, and proximal. Primal agents can be probed given a tentative plan, and they return a corresponding cost of deviation; dual agents can be called with a price, and they return a corresponding favored plan; and proximal agent can consume both a tentative plan and tentative cost, and they return updated values thereof. Most CPP algorithms are built with one type of agent in mind, requiring that all agents belong to the same class. The practical difficulty that we address theoretically in this paper is to devise methodologies that allow for a “mix-and-match” of agents, and are compatible with all types of agent interfaces, so as not to require the preliminary burden of modifying them to fit the same type.
@@ 2.1 Consensus Problem
The consensus problem is one that involves a set of agents $\mathcal{M}$ that each have their own cost function $g_{i},i\in\mathcal{M}$. The goal is to optimize the sum of these functions for a common plan $z$: $\min_{z}\sum_{i\in\mathcal{M}}g_{i}(z)$. In order to leverage the separability of the agents, the problem is re-written to endow each agent with their own plan variable, subject to the constraint that they must all be equal:
$\min_{\{x_{i}\}_{i\in\mathcal{M}},z}$ $\penalty 10000\ \sum_{i\in\mathcal{M}}g_{i}(x_{i})$
$s.t.$ $\penalty 10000\ x_{i}=z,\penalty 10000\ \forall i\in\mathcal{M}.$
@@ 2.2 Agents
In our consensus planning problem, there exist three types of agent interfaces: primal, dual, and proximal. These are the natural interfaces for the three main approaches to the problem: gradient descent, dual ascent, and the augmented Lagrangian method of multipliers, respectively.
@@
$\max_{\{\lambda_{i}\}_{i\in\mathcal{M}}}\left\{\min_{\{x_{i}\}_{i\in\mathcal{M}}}\sum_{i\in\mathcal{M}}g_{i}(x_{i})-\lambda_{i}^{T}(z-x_{i})\right\}.$
We then require from the agents that they be able to consume the dual variables $\lambda_{i}$ and return the solutions to the subproblems in the brackets as $x_{i}=\arg\min_{x}g_{i}(x)-\lambda_{i}^{T}(z-x)$. Note that the dual agent only requires the dual variables since the optimization is over $x$ and the term $\lambda_{i}^{T}z$ is a constant for that problem, which can be omitted.
@@
$\max_{\{\lambda_{i}\}_{i\in\mathcal{M}}}\left\{\min_{\{x_{i}\}_{i\in\mathcal{M}}}\sum_{i\in\mathcal{M}}g_{i}(x_{i})-\lambda_{i}^{T}(z-x_{i})+\frac{\rho}{2}\lVert z-x_{i}\rVert^{2}\right\}.$
@@ 2.3 Previous Work
Consensus planning is often viewed as a variant of distributed optimization, where there are relatively few optimization agents and querying information from each agent is comparatively costly. This variant is in contrast with much of the distributed optimization work where the central concern is that of scale across many agents (ryu2022large, ch.11). Given the requirement of a central coordinator, the setup is similar to that of federated learning.
Most of the distributed optimization algorithms assume the same interface for all agents: for example, a primal interface in the case of distributed (proximal) gradient descent; a dual interface in the case of dual decomposition; and a proximal interface in the case of distributed ADMM. By contrast, this paper allows for any mix of primal, dual, and proximal interfaces.
Heterogeneity of the agents is considered in the context of optimal control, however it usually takes a different meaning. sun2020distributed for example consider coordination across agents that act in either discrete or continuous time.
A closer type of heterogeneity concerns the amount of information and computation that an agent can perform. niu2023fedhybrid consider coordination among agents that can perform either gradient or Newton-type updates, i.e. first or second order agents. In our setting, both would still correspond to what we refer to as primal agents, and although our initial analysis focuses on the first-order formulation of the agents, and we address second order information in the extensions in Section 5.3.
@@ 2.4 Motivating Examples
The different agent interfaces in Table 1 arise naturally from existing system architectures:
Dual interfaces
often correspond to agents that solve pricing-based subproblems (e.g., procurement systems that respond to price signals).
Primal interfaces
correspond to agents that can evaluate costs and gradients but cannot easily solve optimization subproblems (e.g., simulation-based capacity models).
Proximal interfaces
correspond to agents with flexible optimization capabilities (e.g., modern optimization-based planning systems).
This heterogeneity in interfaces reflects the reality of large-scale supply chain systems, where different components were developed at different times with different technological capabilities and constraints. Nonetheless, these systems are usually naturally convex. Buying agents for example typically stem from the modeling of a convex multi-period inventory management problem (zipkin2000foundations), and the dual interface allows for the problem to separate into product-level subproblems. Capacity agents also frequently involve convex costs and linear constraints, leading to convex problems, while transportation or flow agents are also naturally convex (ahuja1988network).
Table 1: Examples of consensus problems at a large online retailer.
@@ 3.1 Derivation
We consider a consensus optimization problem involving a mixed set of agents $i\in\mathcal{M}$ that can be either primal, dual, or proximal. We let $\mathcal{P}$ be the set of primal agents, $\mathcal{D}$ be the set of dual agents, and $\mathcal{X}$ be the set of proximal agents. Each agent’s objective function is given by $g_{i}$, where $i$ denotes their index, and we assume that the functions $g_{i}$ are convex, and in particular $\mu_{i}$-strongly convex for $i\in\mathcal{D}$, and have $\beta_{i}$-Lipschitz continuous gradients for $i\in\mathcal{X}$. The consensus problem to be solved is:
$\min_{\mathbf{x},z}$ $\penalty 10000\ \sum_{i\in\mathcal{M}}g_{i}(x_{i})$ (1)
@@
We first make the different treatments of the agents explicit by separating agents based on their type. Applying partial duality (e.g., (Luenberger2021, ch. 14)) to the dual and proximal agents, and augmenting the latter with the augmented Lagrangian proximal term associated with their corresponding constraint, yields the following equivalent formulation:
$\max_{\lambda_{i},i\in\mathcal{D}\cup\mathcal{X}}\min_{x_{i},z}$ $\penalty 10000\ \sum_{i\in\mathcal{P}}g_{i}(x_{i})+\sum_{i\in\mathcal{D}}g_{i}(x_{i})+\lambda_{i}^{T}(z-x_{i})+\sum_{i\in\mathcal{X}}g_{i}(x_{i})+\lambda_{i}^{T}(z-x_{i})+\frac{\rho_{i}}{2}\lVert z-x_{i}\rVert^{2},$ (2)
@@
Formulation (2) already suggests a direction for the solution of the problem, owing in particular to the similarity in structure between ADMM (boyd2011distributed) and traditional dual ascent methods (Luenberger2021, ch. 14) used to solve dualized consensus problems. Both of these approaches involve iterations comprised of the same 3 steps:
@@
A natural idea is to recast the primal agent as a proximal agent, where its function $g_{i}$ is replaced by some linearized approximation thereof using the gradient information returned by the primal agent. Fortunately, a substantive body of work has demonstrated that the ADMM updates do not need to be exact for the method to converge (he2002new; eckstein1992douglas; yang2011alternating). In particular, the terms of the ADMM updates can be linearized, be it the augmented quadratic term (lin2011linearized; yang2013linearized), or the agent function (ouyang2013stochastic; suzuki2013dual), or both through a more general use of a Bregman divergence term to replace the quadratic penalty, and an additional Bregman divergence term (wang2014bregman). Leveraging these latter results, a simple Bregman-ADMM formulation in the context of the consensus problem has updates of the form:
@@
We will see in Section 3.3 that when all agents are primal, the application of this linearized ADMM yields updates that bear a strong similarity to the ones performed by distributed gradient descent (nedic2009distributed), and to the FedHybrid updates of niu2023fedhybrid.
@@ 3.2 Formulation
We detail in this section the steps of the algorithm following the high level derivation presented in Section 3.1. The agents can be either primal ($i\in\mathcal{P}$), dual ($i\in\mathcal{D}$), or proximal ($i\in\mathcal{X}$). We recall that we consider $\mu_{i}$-strongly convex functions $g_{i},i\in\mathcal{D}$ for dual agents, and $\beta_{i}$-Lipschitz continuous gradients for $g_{i},i\in\mathcal{P}$. We allow for $\mu_{i}=0$ and $L_{i}=+\infty$ for those other functions that are not strongly convex or do not have Lipschitz continuous gradients, for notational simplicity. Each agent is allowed to have their own regularizing parameter/learning rate $\rho_{i}$, although in practice we often use the same one for agents of the same type. We will return to a discussion on the choice of these hyperparameters in Section 5, but since the dual agents essentially perform dual ascent, and as we will confirm in the proof of the convergence results, we impose that $\rho_{i}<\mu_{i},\penalty 10000\ \forall i\in\mathcal{D}$. On the other hand, we recall that we require $L_{i}>\beta_{i}$ for $i\in\mathcal{P}$, and place no restriction on $\rho_{i}>0$ for $i\in\mathcal{X}$. We further assume throughout the paper that the initial prices $\lambda_{i}^{0}$ are such that $\sum_{i\in\mathcal{M}}\lambda_{i}=0$.
@@
$\phi_{i}(x)$ $=\begin{cases}0,&i\in\mathcal{X}\cup\mathcal{D}\\ \frac{L_{i}}{2}\lVert x\rVert^{2}-g_{i}(x),&i\in\mathcal{\mathcal{P}}\end{cases},$ $\tilde{\rho}_{i}$ $=\begin{cases}\rho_{i},&i\in\mathcal{\mathcal{P}\cup\mathcal{X}}\\ 0,&i\in\mathcal{D}.\end{cases}$
Note that, technically, the update of the primal agents is performed by the coordinator. The primal agents are queried with the tentative plan $x_{i}^{k}$ and return $\nabla g_{i}(x_{i}^{k})$, which is then used by the coordinator to yield the updated primal agent plan through (4). Additionally, the consensus update should have a term comprised of a multiple of the sum of the prices, but summing up the price updates, (8) we obtain the optimality condition of the consensus update problem, meaning that the sum of the prices is null after the first iteration (see e.g. boyd2011distributed) [^1] To avoid numerical errors preventing the sum of prices to be null in practice, a simple leveling step can be added after the price updates by subtracting the mean price from each price.. Making the additional assumption that the sum of the initial prices is also null further simplifies the notation. The algorithm is summarized in Algorithm 1, which we name 3-Agent Consensus Planning (3ACP).
@@
Let $\rho_{i}>0,\penalty 10000\ \forall i\in\mathcal{M}$, and $\lambda_{i}^{0}$ such that $\sum_{i\in\mathcal{M}}\lambda_{i}^{0}=0$.
Initialize $x_{i}^{0},\penalty 10000\ \forall i\in\mathcal{M}$ and let $z^{0}:=\frac{1}{\sum_{i\in\mathcal{M}}\rho_{i}}\sum_{i\in\mathcal{M}}\rho_{i}x_{i}^{0}$.
@@ 3.3 Comments
Consider the case when all the agents are primal agents, and assume for simplicity that they use common values $L=L_{i},\forall i$ and $\rho=\rho_{i},\forall i$. The updates take the form of Bregman ADMM (wang2014bregman):
@@
Given that the matrix $W$ is symmetric and doubly stochastic, these updates are almost those of the decentralized gradient descent (nedic2009distributed), except for the presence of dual variables $\lambda^{k}$ in the update. Decentralized gradient descent requires that the step sizes be decreasing in order to converge to the optimal, otherwise it only converges to within a neighborhood of the optimal (nedic2009distributed; yuan2016convergence). When the functions are strongly convex, the consensus plan converges linearly to the optimal until it reaches said neighborhood (yuan2016convergence). By working on both primal and dual variables through the coordinator, the linearized (Bregman) ADMM overcomes the need for decreasing step sizes and converges linearly to the optimal solution.
Another way of expressing the update of the preferred plans when all the agents are primal is as:
$x_{i}^{k+1}$ $=x_{i}^{k}-\frac{1}{L+\rho}\left(\nabla g_{i}(x_{i}^{k})-\lambda_{i}^{k}+\rho(x_{i}^{k}-z^{k})\right),$
which is similar to the gradient-type update of the FedHybrid algorithm (niu2023fedhybrid). A difference being that FedHybrid updates the dual variables using the previous consensus plan, while our approach based on the Bregman ADMM updates the consensus variable before updating the dual variables. Nonetheless, this suggests that FedHybrid could be interpreted as a particular case of Bregman ADMM.
@@
When all the agents are proximal, the algorithm reduces to the traditional consensus ADMM (boyd2011distributed).